Opposite: Beam with pipe cross-section R = 40 mm, r = 38 mm, a = 1 m, structural steel (R. = 235 MPa),

 p = 7.85 kg / dm3, E = 210000 MPa At the end of the pipe there is a mass m hanging on a rope.

 g = 9.81 m / s2

 - 2r 2R

 Ges .: a) Calculate the maximum, taking into account the weight of the pipe

 permissible mass must so that the stress analysis for bending for the pipe is still fulfilled.

 (14 p.)

 b) Give the resulting bending line and the maximum bending deflection for the

 Pipe at mui. The weight of the pipe itself must be taken into account and the coefficients must be calculated before the expressions in brackets.


1
Expert's answer
2021-02-11T10:41:48-0500

Given question is incomplete, we can proceed this type of question by using the following formula-

"I=\frac{\pi (R^4-r^4)}{64}"


"\Rightarrow I = \frac{\pi \times (40^4-38^4)}{64}\times 10^{-12}"

Now, for the calculation of the bending stress

"\frac{M_b}{I}=\frac{\sigma_b}{y}"

here "M_b=\frac{wl}{4}"

now we will substituting the values in the above and calculate the required values.